OpenStudy (anonymous):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1 over 3π (x - 7)2 (x - 3) What does the factor 1 over 3(x - 7) represent?

OpenStudy (anonymous):

The total surface area of the tank One third of the radius of the tank One third of the height of the tank The total area of the base of the tank

OpenStudy (anonymous):

:l

OpenStudy (anonymous):

Oh and also, can someone answer this: 4(-2) - 3 - 2(-2) ,,, Is it -7?

OpenStudy (anonymous):

@peachpi @xapproachesinfinity help?

OpenStudy (xapproachesinfinity):

the volume of a cone is 1/3 the volume of cylinder

OpenStudy (xapproachesinfinity):

what is the volume of a cylinder

OpenStudy (anonymous):

I dunno

OpenStudy (xapproachesinfinity):

volume = pi (radius squared) (height)

OpenStudy (xapproachesinfinity):

volume of a cone =1/3 pi (radius squared) (height)

OpenStudy (xapproachesinfinity):

we have the height as x-3

OpenStudy (xapproachesinfinity):

4 less than the height would be x-3-4=x-7 and that's 5he radius

OpenStudy (anonymous):

so that means... im sorry I haven't done this in a super long time